# Yuri Balashov

### v1

@@ 5.1 In-Context Translation
Few-shot translation mentioned above (Section 4.1) is a remarkable example of in-context learning in LLMs (Brown et al., 2020; Dong et al., 2024). The “unreasonable effectiveness” (Garcia et al., 2023) of in-context translation and its ability to override semantically inadequate prompt instructions (Zhu et al., 2024) raises questions about its underlying computational mechanisms.
Sia et al. (2024) approached these questions by asking at what stage in the computation LLMs recognize and execute the translation task in a few-shot mode. By using a technique called “layer-from context masking,” the authors selectively removed the model’s ability to attend to prompt instructions and examples from specific layers onward. This method allowed them to locate a “task recognition point”—around layer 14 of 32 in Llama 7B, for instance—where the model has internalized the task and no longer needs to consult the context to generate accurate translations. Before this point, masking the context significantly degraded performance; after it, performance remained stable even when the context was masked. This suggests a three-phase process: an initial stage with minimal context influence, a middle stage of active task location, and a final stage where execution proceeds independently of the prompt.
@@ 5.2 Cross-Lingual Representation Alignment
An intriguing earlier finding by Wen-Yi and Mimno (2023) was that some language models cluster semantically similar words from different languages together at the very first, often ignored, embedding layer. Specifically, the authors found that, without explicit bilingual training, the token embeddings of mT5-XL (Xue et al., 2021) often align across languages for concepts; for example, the vectors for når (“when,” Danish), cuando (“when,” Spanish), quando (“when,” Portuguese), \<k/s> (“when,” Hebrew), and quand (“when,” French) are all close neighbors. A similar phenomenon was earlier observed in models like mBERT (Devlin et al., 2019), but not in the input embedding layers.
@@
What about pristine decoder-only models? Are they also capable of learning cross-lingual representation alignment and cross-lingual transfer during pre-training? Hua et al. (2024) approach these questions with a controlled training of GPT-2 models on a multilingual extension of a toy synthetic language that was earlier used to model legal moves in the Othello board game (Li et al., 2024). While limited both in scope and scale, their findings suggest that relatively small decoder-only models do not learn a common representation across multiple artificial languages when the models are exposed to them separately. However, adding “anchor tokens”—lexical items that are shared across languages—to the training sequences improves cross-lingual representation alignment quite considerably. As the authors note, mOthello languages with their simplistic grammar and only 180 tokens are a far cry from natural languages. Yet the catalyzing role of “anchor tokens” in “Global” (Section 1) multilingual representation learning is notable and consistent with the role of shared subword vocabularies and “seed lexicons” earlier studied in other settings, such as unsupervised MT (Artetxe et al., 2018; Søgaard et al., 2018; Conneau et al., 2020a).
@@ 5.3 Do LLMs Pivot Through English on Their Own?
Already mentioned in Section 4.3, pivoting through an intermediate language (most often English) has been widely exploited in the MT industry for a long time. Recent user studies have extended this approach to LLMs by explicitly asking the models to go through two stages, and sometimes found improved accuracy (see, e.g., Jiao et al. 2023b).
Interestingly, evidence suggests LLMs themselves might internally be doing something like this. Wendler et al. (2024) provided insight by asking: do multilingual LLMs use English as an internal pivot language? They prompted Llama-2 models to translate a single French word ‘fleur’ to Chinese with a few-shot prompt (where ‘{CJK*}UTF8bsmi中文’ means “Chinese”):
Francais: "vertu" - {CJK*}UTF8bsmi中文: "{CJK*}UTF8bsmi德" Francais: "siege" - {CJK*}UTF8bsmi中文: "{CJK*}UTF8bsmi座 Francais: "neige" - {CJK*}UTF8bsmi中文: "{CJK*}UTF8bsmi雪 Francais: "montagne" - {CJK*}UTF8bsmi中文: "{CJK*}UTF8bsmi山 Francais: "fleur" - {CJK*}UTF8bsmi中文: "
and applied a technique called logit lens (Nostalgebraist, 2020) to unembed models’ intermediate layers early to see what they encode. They found that the model’s representation of the French input drifts closer to English semantic space before settling into the target language space (i.e. Chinese) (Figure 2).
Figure 2: Llama-2 7B translates ‘fleur’ to ‘{CJK*}UTF8bsmi花’ by seemingly pivoting through ‘flower’ in intermediate layers. Figure 1 in Wendler et al. (2024).
In essence, an LLM might internally translate everything to a kind of “English-centric interlingua” as a steppingstone. If so, then explicitly pivoting via English might resonate with the model’s workings. Indeed, some prompting approaches encourage the model to think aloud in English before producing the final translation in the target language. And more recent “reasoning” models sometimes do it on their own (Liu et al., 2025; Chen et al., 2025), using English as a kind of “metalanguage.” [^8] Unfortunately, this interesting metalinguistic phenomenon lies outside the scope of the paper. Could it form the basis of another distinct mode of translation as ”extended inference” analogous to similar operations sometimes performed by human translators when they encounter difficult cases? This question is definitely worth exploring. Thanks to Raphaël Millière for drawing my attention to it.
@@ 5.4 Language-Agnostic Representation?
But do LLMs actually “think” in English when performing multilingual tasks (Schut et al., 2025)? Or do they “think” in a truly interlingual way, with English emerging in the intermediate layers simply because it happens to dominate the training data? Wendler et al.’s early decoding study (Wendler et al., 2024) did not allow them to address this question head-on. Dumas et al. (2025) take up this challenge with a more nuanced mechanistic interpretability approach based on activation patching. They designed simple translation prompts in different languages to serve as two contexts: a “source prompt” (translating a word from, say, German to Italian) and a “target prompt” (translating a different word from, say, French to Chinese). The authors ran Llama-2 7B on both prompts and then swapped part of the internal state between them mid-process. More precisely, at a chosen transformer layer, they took the hidden latent representation of the last token from the source prompt’s forward pass and injected it into the target prompt’s forward pass at the same layer. By observing how this surgical insertion alters the model’s next-word prediction, they can tell what information that latent was carrying (Figure 3).
The outcome was quite striking: different layers separated the handling of language and meaning. When Dumas et al. (2025) patched at early layers (closer to the input), Llama-2 still produced the correct word in the target language, as if nothing had changed. But patching at a middle layer caused a curious swap: the model generated the correct concept but in the wrong language (specifically, it used the source prompt’s language). And patching at a late layer flipped both aspects—the concept as well as the language, effectively following the source prompt’s content entirely. This layered behavior reveals that Llama-2 selects the output language earlier in the network, and determines the actual concept to translate later, in deeper layers. In other words, the model first decides “I need to respond in Chinese” (target language) before it figures out “what is a lemon?” in the context of translation.
[…]
@@
The authors note the limitations of their study which operates with single word prompts and very simple unambiguous concepts, explicitly acknowledging that richer lexical items, multi-word expressions, sentential context, discourse phenomena and language-specific idioms remain outside the scope of this approach. It is unclear whether the demonstrated disentanglement of language and meaning carries over to full-sentence translation. All the reported experiments use open-source models (in addition to Llama-2 7B, the authors ran experiments with Llama-3 8B, Mistral 7B, Qwen 7B, Aya 8B, Gemma 2B). While the authors replicated some of their findings on a 70B Llama, they obviously could not test frontier-scale, instruction-tuned or RLHF-aligned systems. Differences in the model size, training mix, and alignment objectives may alter how larger “consumer-grade” LLMs partition language and meaning. Finally, while activation patching is a powerful tool it may introduce artifacts and does not, by itself, guarantee that the model would naturally traverse the identified internal states during ordinary inference. But despite these limitations, Dumas et al.’s work contributes important insights to the ongoing debate about the nature of multilingual representation in LLMs.
A further step in this direction was recently undertaken by Anthropic’s transformer circuits team (Lindsey et al., 2025) who used a newly developed tool called attribution graphs to trace how certain features (interpretable patterns of activation in the network) interact to produce the model’s output.
@@
In a bit more detail, the tool was put to work in a simple translation-like task: the prompt “The opposite of ‘small’ is …” posed in three different languages (English, French, and Chinese). Despite the surface language differences, all three prompts have the same meaning, and their correct completions are semantically equivalent: ‘big’, ‘grand’, and ‘{CJK*}UTF8bsmi大’. The question is: does the model (Claude 3.5 Haiku) respond to these prompts using one language-neutral reasoning process, or does it handle each language separately? The attribution graphs reveal a striking answer: the model uses very similar internal circuits for all three languages combining shared cross-lingual features with language-specific ones. In each case, Claude first activates a set of language-independent features that recognize the task of finding the antonym of ‘small’. These include an abstract concept of “smallness” and an “opposite-of” operation feature. Once this conceptual step finds the idea of big (the opposite of small), the model then engages language-specific features to express that idea in the appropriate tongue. In other words, Claude appears to “think” in a language-neutral way about the meaning (identifying the “opposite of small” as the concept big), and only then translate that concept into the word forms ‘big’, ‘grand’, or ‘{CJK*}UTF8bsmi大’ depending on whether the context is English, French, or Chinese. There are features that act as language markers; for example, an open-quote-in-French feature or a Chinese-context feature, which ensure the answer is framed in the correct language with the right quotation marks. But the core semantic features (for small and its opposite) are multilingual: the model has learned a single abstract notion of small that spans languages, rather than distinct unrelated notions for each language. The authors describe the overall computation as having three factorable parts: Operation (finding an antonym), Operand (the concept of small), and Language (which language’s vocabulary to use for the answer). This decomposition suggests that the model might indeed be employing something like an internal “interlingua” for meaning, combined with a separate layer for language-specific expression.
@@
The results confirmed the expectation that Claude’s internal processing has a “modular” structure: one module figures out the task (antonym vs. synonym), another handles the core meaning of the words involved, and a third handles the language context. Each can be perturbed on its own. This level of interpretability is quite remarkable: it is as if one opened the model’s “brain” and identified a sub-network for “opposite meaning,” a sub-network for “the concept of small,” and a sub-network for “speaking French,” and demonstrated each one’s role by poking it. This offers a fascinating parallel to the idea that meanings are not intrinsically tied to words: the model generates the appropriate word when needed, implying it has an independent grasp of the meaning itself. Furthermore, the model’s ability to swap out Operation and Operand independently points to a form of compositional representation inside the neural network. The classical hallmark of compositionality in linguistic semantics is the requirement that “the meaning of an expression [be] a function of the meanings of its parts and of the way they are syntactically combined” (Partee, 1984, 281). Here we see it at work: the model’s internal process for “the opposite of X” can be decomposed into a part that understands the relation “opposite-of” and a part that provides the content “X,” and these parts can be intervened on separately. The “antonym-finding” operation acts almost like a built-in function that can be applied to different inputs (‘small’ or ‘hot’ or others) to yield different outputs. Importantly, the concept of small is represented in a way that is not entangled with the fact that it was asked in French or English; it is a “portable” piece of information.
@@
Figure 4: Three targeted interventions on Claude’s multilingual circuit (Lindsey et al. 2025, Figure 18). Left: Operation Swap—replacing the antonym operation features with synonym features (while the input still asks for “opposite”) causes the model to output ‘little’ (a synonym of ‘small’) instead of the antonym ‘big’. This shows that toggling the internal operation from “find opposite” to “find similar” yields a corresponding change in output. Center: Operand Swap—replacing the features encoding the concept small with those for hot causes the model to respond with ‘cold’ (the antonym of ‘hot’). The prompt remains “The opposite of ‘small’ is…”, but internally the model now treats ‘small’ as if it were ‘hot’, demonstrating that the content of the adjective can be swapped independently of the prompt’s wording. Right: Language Swap—replacing English-language context features with Chinese ones makes the model output ‘{CJK*}UTF8bsmi大’ (Chinese for ‘big’) for the English prompt. The model still found the concept big as the opposite of small but expressed it in a different language. In each case, only one aspect of the computation is changed (function, content, or language), and the model’s behavior changes in the predicted way, confirming that these internal features indeed govern those distinct components of the translation task.
@@
On a cautionary note, Anthropic’s team also observed that English seems to be the model’s “default” language in this internal process; which is hardly surprising. They describe English as “mechanistically privileged,” meaning that, in the model’s circuits, English has a built-in advantage or priority. How does this show up mechanistically? It turns out that those abstract “say big” features (the ones that correspond to expressing the concept of big) have a stronger direct influence on producing the English word ‘big’ than the equivalents in French or Chinese. In Claude’s neural wiring, the path from the concept of big to the English token ‘big’ is somewhat more straightforward and higher-weighted than the path to ‘grand’ or ‘{CJK*}UTF8bsmi大’. For non-English outputs, the model relies a bit more on the extra “language context” features to get to the right word. Another subtle observation was that some English-specific features engage in what the authors call a “double inhibitory effect”: essentially, certain English features suppress competing features that would otherwise produce the English word when the context is not English, thus allowing the non-English word to come out. This complicated dance of inhibition and activation paints a picture in which the model retains a universal semantic representation (of big) that is biased to map to English unless additional mechanisms redirect it. In more philosophical terms, Claude 3.5 has a kind of universal “mentalese” for concepts (Fodor, 1975), but it is skewed towards outputting English by default. Only when the model clearly detects a non-English context (quotes, characters, etc.) do these language-specific signals intervene to translate that internal thought into, say, French. This finding echoes the tension noted in the recent literature: some prior work had found evidence of a shared multilingual representation (Zhang et al., 2024), while others noted an English-centric bias (Schut et al., 2025). Here, both appear to be true: the representations are genuinely multilingual, and English is a privileged default internally. One could say the model has a universal “mental dictionary” but with English entries written in bold.

### v3

@@ 5.1 In-Context Translation
Few-shot translation mentioned above (Section 4.1) is a remarkable example of in-context learning in LLMs (brown_language_2020; dong_survey_2024). The “unreasonable effectiveness” (garcia_unreasonable_2023) of in-context translation and its ability to override semantically inadequate prompt instructions (zhu_multilingual_2024) raise questions about its underlying computational mechanisms.
sia_where_2024 approached these questions by asking at what stage in the computation LLMs recognize and execute the translation task in a few-shot mode. By using a technique called “layer-from context masking,” the authors selectively removed the model’s ability to attend to prompt instructions and examples from specific layers onward. This method allowed them to locate a “task recognition point”—around layer 14 of 32 in Llama 7B, for instance—where the model has internalized the task and no longer needs to consult the context to generate accurate translations. Before this point, masking the context significantly degraded performance; after it, performance remained stable even when the context was masked. This suggests a three-phase process: an initial stage with minimal context influence, a middle stage of active task location, and a final stage where execution proceeds independently of the prompt.
@@ 5.2 Cross-Lingual Representation Alignment
An intriguing earlier finding by wen-yi_hyperpolyglot_2023 was that some language models cluster semantically similar words from different languages together at the very first, often ignored, embedding layer. Specifically, the authors found that, without explicit bilingual training, the token embeddings of mT5-XL (xue_mt5_2021) often align across languages for concepts; for example, the vectors for når (“when,” Danish), cuando (“when,” Spanish), quando (“when,” Portuguese), k/s! (“when,” Hebrew), and quand (“when,” French) are all close neighbors. A similar phenomenon was earlier observed in models like mBERT (devlin_bert_2019), but not in the input embedding layers.
@@
What about pristine decoder-only models? Are they also capable of learning cross-lingual representation alignment and cross-lingual transfer during pre-training? hua_mothello_2024 approach these questions with a controlled training of GPT-2 models on a multilingual extension of a toy synthetic language that was earlier used to model legal moves in the Othello board game (li_emergent_2024). While limited both in scope and scale, their findings suggest that relatively small decoder-only models do not learn a common representation across multiple artificial languages when the models are exposed to them separately. However, adding “anchor tokens”—lexical items that are shared across languages—to the training sequences improves cross-lingual representation alignment quite considerably. As the authors note, mOthello languages with their simplistic grammar and only 180 tokens are a far cry from natural languages. Yet the catalyzing role of “anchor tokens” in “Global” (Section 1) multilingual representation learning is notable and consistent with the role of shared subword vocabularies and “seed lexicons” earlier studied in other settings, such as unsupervised MT (artetxe_unsupervised_2018; sogaard_limitations_2018; conneau_unsupervised_2020) and multilingual word embeddings more generally (koehn_neural_2020, Ch. 12).
Experiments with toy models, synthetic data, and narrowly scoped tasks offer a distinctive advantage: they give researchers fine-grained control over training dynamics. In a recent study, blum2025rosettastoneunificationforces trained tiny 2M-parameter transformers from scratch on synthetic multilingual corpora and uncovered an early learning phase during which models either unify or segregate the same facts across languages. Only the unification regime yields robust cross-lingual transfer. The authors show that the strength of unification closely tracks cross-lingual factual accuracy and can be causally modulated through dataset design. In particular, when language identity is easy to infer and predictive of simple labels, models drift toward language-siloed representations; by contrast, balancing the data and attenuating language-specific surface cues encourages cross-lingual unification.
@@ 5.3 Do LLMs Pivot Through English on Their Own?
Already mentioned in Section 4.3, pivoting through an intermediate language (most often English) has been widely exploited in the MT industry for a long time. Recent user studies have extended this approach to LLMs by explicitly asking the models to go through two stages, and sometimes found improved accuracy.
Interestingly, evidence suggests LLMs themselves might internally be doing something like this. wendler_llamas_2024 provided insight by asking: do multilingual LLMs use English as an internal pivot language? They prompted Llama-2 models to translate a single French word ‘fleur’ to Chinese with a few-shot prompt (where ‘中文’ means “Chinese”):
Francais: "vertu" - 中文: "德" Francais: "siege" - 中文: "座" Francais: "neige" - 中文: "雪" Francais: "montagne" - 中文: "山" Francais: "fleur" - 中文: "
and applied a technique called logit lens (nostalgebraist_interpreting_2020) to unembed models’ intermediate layers early to see what they encode. They found that the model’s representation of the French input drifts closer to English semantic space before settling into the target language space (i.e. Chinese) (Figure 4).
Figure 4: Llama-2 7B translates ‘fleur’ to ‘花’ by seemingly pivoting through ‘flower’ in intermediate layers. Figure 1 in wendler_llamas_2024.
In essence, an LLM might internally translate everything to a kind of “English-centric interlingua” as a steppingstone. If so, then explicitly pivoting via English might resonate with the model’s inner workings. Indeed, some prompting approaches encourage the model to think aloud in English before producing the final translation in the target language. And more recent “reasoning” models sometimes do it on their own (liu_new_2025; chen_evaluating_2025), using English as a kind of “metalanguage.” [^7] Unfortunately, this interesting metalinguistic phenomenon lies outside the scope of the paper. Could it form the basis of another distinct mode of translation as ”extended inference” analogous to similar operations sometimes performed by human translators when they encounter difficult cases? This question is definitely worth exploring. Thanks to Raphaël Millière for drawing my attention to it.
@@ 5.4 Language-Agnostic Representation?
But do LLMs actually “think” in English when performing multilingual tasks (schut_multilingual_2025)? Or do they “think” in a truly interlingual way, with English emerging in the intermediate layers simply because it happens to dominate the training data (wang-etal-2025-lost-multilinguality)? Wendler et al.’s early decoding study (wendler_llamas_2024; lim2025languagespecificlatentprocesshinders) did not allow them to address this question head-on. dumas_separating_2025 take up this challenge with a more nuanced mechanistic interpretability approach based on activation patching. They designed simple translation prompts in different languages to serve as two contexts: a “source prompt” (translating a word from, say, German to Italian) and a “target prompt” (translating a different word from, say, French to Chinese). The authors ran Llama-2 7B on both prompts and then swapped part of the internal state between them mid-process. More precisely, at a chosen transformer layer, they took the hidden latent representation of the last token from the source prompt’s forward pass and injected it into the target prompt’s forward pass at the same layer. By observing how this surgical insertion alters the model’s next-word prediction, they can tell what information that latent was carrying (Figure 5).
Figure 5: Figure 1 in dumas_separating_2025. A hidden state from the source prompt (red, left) at a certain layer is inserted into the forward pass of the target prompt (blue, right) at the corresponding layer. The source prompt here provides an example translation (German ‘Dorf’ $\rightarrow$ Italian ‘villaggio’) and then asks to translate ‘Buch’ (book) to Italian, while the target prompt asks to translate French ‘citron’ (lemon) to Chinese. By patching the red latent into the blue context at different layers, the researchers observed changes in the predicted translation. Early-layer patching yields the correct target concept in the target language (e.g. ‘柠檬’ for lemon in Chinese), mid-layer patching yields the correct concept but in the wrong language (‘limone’ = lemon in Italian), and late-layer patching yields the wrong concept in the source language (‘libro’ = book in Italian).
The outcome was quite striking: different layers separated the handling of language and meaning. When dumas_separating_2025 patched at early layers (closer to the input), Llama-2 still produced the correct word in the target language, as if nothing had changed. But patching at a middle layer caused a curious swap: the model generated the correct concept but in the wrong language (specifically, it used the source prompt’s language). And patching at a late layer flipped both aspects—the concept as well as the language, effectively following the source prompt’s content entirely. This layered behavior reveals that Llama-2 selects the output language earlier in the network, and determines the actual concept to translate later, in deeper layers. In other words, the model first decides “I need to respond in Chinese” (target language) before it figures out “what is a lemon?” in the context of translation.
@@
The authors note the limitations of their study which operates with single word prompts and very simple unambiguous concepts, explicitly acknowledging that richer lexical items, multi‑word expressions, sentential context, discourse phenomena and language‑specific idioms remain outside the scope of this approach. It is unclear whether the demonstrated disentanglement of language and meaning carries over to full-sentence translation. All the reported experiments use open-source models (in addition to Llama‑2 7B, the authors ran experiments with Llama‑3 8B, Mistral 7B, Qwen 7B, Aya 8B, Gemma 2B). While the authors replicated some of their findings on a 70B Llama, they obviously could not test frontier‑scale, instruction‑tuned or RLHF‑aligned systems. Differences in the model size, training mix, and alignment objectives may alter how larger “consumer‑grade” LLMs partition language and meaning. Finally, while activation patching is a powerful tool it may introduce artifacts and does not, by itself, guarantee that the model would naturally traverse the identified internal states during ordinary inference. But despite these limitations, Dumas et al.’s work contributes important insights to the ongoing debate about the nature of multilingual representation in LLMs.
A further step in this direction was recently undertaken by Anthropic’s transformer circuits team (lindsey_biology_2025) who used a newly developed tool called attribution graphs to trace how certain features (interpretable patterns of activation in the network) interact to produce the model’s output.
@@
In a bit more detail, the tool was put to work in a simple translation-like task: the prompt “The opposite of ‘small’ is …” posed in three different languages (English, French, and Chinese). Despite the surface language differences, all three prompts have the same meaning, and their correct completions are semantically equivalent: ‘big’, ‘grand’, and ‘大’. The question is: does the model (Claude 3.5 Haiku) respond to these prompts using one language-neutral reasoning process, or does it handle each language separately? The attribution graphs reveal a striking answer: the model uses very similar internal circuits for all three languages combining shared cross-lingual features with language-specific ones. In each case, Claude first activates a set of language-independent features that recognize the task of finding the antonym of ‘small’. These include an abstract concept of “smallness” and an “opposite-of” operation feature. Once this conceptual step finds the idea of big (the opposite of small), the model then engages language-specific features to express that idea in the appropriate tongue. In other words, Claude appears to “think” in a language-neutral way about the meaning (identifying the “opposite of small” as the concept big), and only then translate that concept into the word forms ‘big’, ‘grand’, or ‘大’ depending on whether the context is English, French, or Chinese. There are features that act as language markers; for example, an open-quote-in-French feature or a Chinese-context feature, which ensure the answer is framed in the correct language with the right quotation marks. But the core semantic features (for small and its opposite) are multilingual: the model has learned a single abstract notion of small that spans languages, rather than distinct unrelated notions for each language. The authors describe the overall computation as having three factorable parts: Operation (finding an antonym), Operand (the concept of small), and Language (which language’s vocabulary to use for the answer). This decomposition suggests that the model might indeed be employing something like an internal “interlingua” for meaning, combined with a separate layer for language-specific expression.
@@
Figure 6: Three targeted interventions on Claude’s multilingual circuit (lindsey_biology_2025, Figure 18). Left: Operation Swap—replacing the antonym operation features with synonym features (while the input still asks for “opposite”) causes the model to output ‘little’ (a synonym of ‘small’) instead of the antonym ‘big’. This shows that toggling the internal operation from “find opposite” to “find similar” yields a corresponding change in output. Center: Operand Swap—replacing the features encoding the concept small with those for hot causes the model to respond with ‘cold’ (the antonym of ‘hot’). The prompt remains “The opposite of ‘small’ is…”, but internally the model now treats ‘small’ as if it were ‘hot’, demonstrating that the content of the adjective can be swapped independently of the prompt’s wording. Right: Language Swap—replacing English-language context features with Chinese ones makes the model output ‘大’ (Chinese for ‘big’) for the English prompt. The model still found the concept big as the opposite of small but expressed it in a different language. In each case, only one aspect of the computation is changed (function, content, or language), and the model’s behavior changes in the predicted way, confirming that these internal features indeed govern those distinct components of the translation task.
The results confirmed the expectation that Claude’s internal processing has a “modular” structure: one module figures out the task (antonym vs. synonym), another handles the core meaning of the words involved, and a third handles the language context. Each can be perturbed on its own. This level of interpretability is quite remarkable: it is as if one opened the model’s “brain” and identified a sub-network for “opposite meaning,” a sub-network for “the concept of small,” and a sub-network for “speaking French,” and demonstrated each one’s role by poking it. This offers a fascinating parallel to the idea that meanings are not intrinsically tied to words: the model generates the appropriate word when needed, implying it has an independent grasp of the meaning itself. Furthermore, the model’s ability to swap out Operation and Operand independently points to a form of compositional representation inside the neural network. The classical hallmark of compositionality in linguistic semantics is the requirement that “the meaning of an expression [be] a function of the meanings of its parts and of the way they are syntactically combined” (partee_barbara_compositionality_1984, 281). Here we see it at work: the model’s internal process for “the opposite of X” can be decomposed into a part that understands the relation “opposite-of” and a part that provides the content “X,” and these parts can be intervened on separately. The “antonym-finding” operation acts almost like a built-in function that can be applied to different inputs (‘small’ or ‘hot’ or others) to yield different outputs. Importantly, the concept of small is represented in a way that is not entangled with the fact that it was asked in French or English; it is a “portable” piece of information.
@@
On a cautionary note, Anthropic’s team also observed that English seems to be the model’s “default” language in this internal process; which is hardly surprising. They describe English as “mechanistically privileged,” meaning that, in the model’s circuits, English has a built-in advantage or priority. How does this show up mechanistically? It turns out that those abstract “say big” features (the ones that correspond to expressing the concept of big) have a stronger direct influence on producing the English word ‘big’ than the equivalents in French or Chinese. In Claude’s neural wiring, the path from the concept of big to the English token ‘big’ is somewhat more straightforward and higher-weighted than the path to ‘grand’ or ‘大’. For non-English outputs, the model relies a bit more on the extra “language context” features to get to the right word. Another subtle observation was that some English-specific features engage in what the authors call a “double inhibitory effect”: essentially, certain English features suppress competing features that would otherwise produce the English word when the context is not English, thus allowing the non-English word to come out. This complicated dance of inhibition and activation paints a picture in which the model retains a universal semantic representation (of big) that is biased to map to English unless additional mechanisms redirect it. In more philosophical terms, Claude 3.5 has a kind of universal “mentalese” for concepts (Fodor1975-FODTLO), but it is skewed towards outputting English by default. Only when the model clearly detects a non-English context (quotes, characters, etc.) do these language-specific signals intervene to translate that internal thought into, say, French. This finding echoes the tension noted in the recent literature: some prior work had found evidence of a shared multilingual representation (zhang_same_2024), while others noted an English-centric bias (schut_multilingual_2025). Here, both appear to be true: the representations are genuinely multilingual, and English is a privileged default internally. One could say the model has a universal “mental dictionary” but with English entries written in bold.